# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ 15.5 The Exact One-Pole Solution and Minkowski Transition
$\Delta X_{0}=\int_{0}^{2\pi}\sqrt{2}R\sin\left((1-a)\theta\right)d\theta=\frac{2\sqrt{2}R\sin^{2}(\pi a)}{1-a}$ (15.7)
We balance the string tension against the boundary energy by minimizing the combined classical action $S(R)=S_{M,Area}-E\Delta X_{0}$ with respect to the geometric scale $R$. Since the Area is purely quadratic $(AR^{2})$ and the boundary coupling is strictly linear $(BR)$, the minimization of this quadratic function trivially yields $B^{2}/(4A)$ by the virial theorem. This dynamically eliminates the dimensionful size $R$ in one step, leaving a purely intensive, remarkably pristine geometric boundary potential:
$S_{geom}(a)=-\frac{E^{2}}{\pi\sigma}F(a),\qquad\text{where}\qquad F(a)=\frac{2\sin^{4}(\pi a)}{1-a}$ (15.8)
This establishes the exact, scale-independent geometric potential that governs the twistor branch point degree $a$ at the WKB saddle point.
15.6 Quark Spin Monodromy, Trace Splitting, and the Unified Spectrum
The final component of the path integral is the path-ordered Dirac trace $T$ along the quark boundary. The global boundary holonomy matrix acting on the twistor space is $U=\text{diag}(\exp\left(-2\pi\mathrm{i}a\right),1)$.
1. The $\rho$ Trajectory (Vector Current)
With an extra $\gamma_{\nu}$ projector inserted as the vector quark current, the twistor dyadic projector $P\propto\lambda\mu^{\dagger}$ acts as a strict rank-1 projector, $\left(\begin{smallmatrix}1&1\\ 0&0\end{smallmatrix}\right)$. It annihilates the trivial constant mode of the un-twisted right-chiral twistor, leaving exactly:
$T_{\rho}\propto\text{tr}(UP)$ $=\text{tr}\begin{pmatrix}\exp\left(-2\pi\mathrm{i}a\right)&\exp\left(-2\pi\mathrm{i}a\right)\\ 0&0\end{pmatrix}$
$=\exp\left(-2\pi\mathrm{i}a\right)$ (15.9)
This corresponding pure topological phase factor $\exp\left(-2\pi\mathrm{i}a\right)$ adds directly to the rotation weight $\exp\left(4\pi\mathrm{i}Ja\right)$, shifting the angular momentum exactly by $\Delta J=+1/2$.
2. The $\pi$ and $K$ Trajectories (Pseudoscalar Current)
With a $\gamma_{5}$ inserted for the pseudoscalar and scalar trajectories, the trace evaluates the sum of the modes:
$T_{\pi,K}=\text{tr}(U)=1+\exp\left(-2\pi\mathrm{i}a\right)$ (15.10)
Because the path integral weight is linear in $T$, the amplitude rigorously splits into a sum of two independent path integrals:
  • 1.
    The Light Integral ($T_{1}=1$): Yields a phase $\exp\left(0\right)$. The spin trajectory is completely unshifted. This generates the ultra-light pseudo-Goldstone trajectory (the $\pi$ and $K$).
  • 2.
    The Heavy Integral ($T_{2}=\exp\left(-2\pi\mathrm{i}a\right)$): Yields the phase $\exp\left(-2\pi\mathrm{i}a\right)$. It acts identically to the $\rho$ projector, generating a heavy vector trajectory rigidly shifted by $\Delta J=1/2$ to $J=1$.
15.7 Asymptotic Linearity and the Chew-Frautschi Plot
In the macroscopic limit ($a\to 1/2$), the parametric geometric equations generate an exact simple pole, aggressively projecting the system into asymptotic Regge linearity. The theoretical intercepts emerge natively from the Liouville anomaly $\frac{\pi a^{2}}{12}$ without a single adjustable parameter:
  • 1.
    Pseudoscalar ($q=0$): $\alpha_{\pi}(0)=-1/96$
  • 2.
    Vector (q=1q=-1): $\alpha_{\rho}(0)=-1/96+1/2=47/96$
Figure 15: Exact Twistor String Trajectories $J=\alpha(m^{2})$. The plot demonstrates the remarkable geometric linearity of the transcendental equations. The $\pi$ and $K$ mesons reside on the unshifted trajectory (with the Kaon explicitly shifted by the perturbative $SU(3)$ flavor breaking $\Delta m_{s}^{2}$), while the vector $\rho$ family is natively shifted upward by $\Delta J=1/2$ due to the rank-1 twistor projection. Unweighted least-squares fits of $\sigma$ map the universal curves through the experimental PDG data.
The physical branch is covered by $a_{\mathrm{min}}<a<1/2$. Here $a_{\mathrm{min}}$ is a solution of the transcendental equation
$4\pi\cot(\pi a)$ $=\frac{1-2a}{a(1-a)}+\frac{1}{a}$
$\implies a_{\mathrm{min}}$ $\approx 0.426836\dots$ (15.11)
Such a narrow range of the parameter $a$ explains the approximate linearity of the Regge trajectories. We computed the asymptotic expansion (G.11) of the Regge trajectory at large $s=m^{2}/\sigma$. The linear and constant terms dominate, and the higher-order $1/s$ corrections are numerically so small that at $J>1$ the deviations from linearity are invisible. But this is, of course, a matter of principle to have a correct formula based on an exact analytic solution instead of heuristic models.
An interesting property of this analytic formula, discussed in G.2, is that these trajectories do not extend to negative spin or to negative $m^{2}$ (tachyons). Instead, they terminate at a geometric singularity characterized by the simultaneous vanishing of the derivatives of our parametric equations, $dJ/da=dm^{2}/da=0$. This singularity occurs at a certain critical value of $a=a_{c}<a_{\mathrm{min}}$ that is safely below the end of the physical region $a_{\mathrm{min}}<a<1/2$. Therefore, this singularity does not manifest in the physical region, where the trajectories appear virtually linear with the corresponding asymptotic intercepts 1/96-1/96 and $47/96$. Thus, this geometric singularity serves as a defender against tachyons. The technical details can be found in G.2.
As shown in Figure 15, the analytic solution maps well onto the experimental data. What is truly remarkable is the numerical linearity of these trajectories given the highly transcendental nature of the geometric equations. The complex trigonometric terms act as deep-IR regulators, providing the correct chiral symmetry breaking behavior for the anomalously light pion. Yet, driven by the exact geometric pole at $a\to 1/2$, the macroscopic strings violently straighten out past $J\approx 1$, validating the strict Regge limit while rigorously housing the microscopic quantum anomalies.
The unweighted least-squares fits yield effective string tensions $\sigma_{\rho}\approx 0.235$, $\sigma_{\pi}\approx 0.281$, and $\sigma_{K}\approx 0.320\text{ GeV}^{2}$. The value for the natural-parity $\rho$ trajectory recovers the canonical macroscopic Regge slope $\alpha^{\prime}=2/(\pi^{2}\sigma_{\rho})\approx 0.86\text{ GeV}^{-2}$, while the slightly higher effective tensions for the $\pi$ and $K$ trajectories physically reflect the dynamical stiffening due to $N_{c}=3$ chiral symmetry breaking and the endpoint inertia of the massive strange quark, respectively.
15.8 Special case of rational branch points
[…]
These rational values of $a$ would not correspond, in general, to integer or half-integer values of the angular momentum $J+q/2$ in (G.9) . In the limit $a\to 1/2$, the angular momentum tends to $\infty$ in this equation. The solutions $a_{l}(J,q)$ of the equation $g(a)=J+q/2$ are, in general, transcendental, so that the physical spectrum is not affected by this geometric extension. However, this phenomenon may be interesting from a purely mathematical point of view, as a number-theoretic exception to the Exact WKB quantization. In the H we rigorously prove that this is not the case for the physical branch , covered by $a_{\textbf{min}}<a<1/2$, where $a_{\textbf{min}}\approx 0.426836$... was introduced above.
16 Planar QCD as a Geometric theory
The dynamical generation of S-matrix mass poles through the degeneracy of the twistor-string effective action is not an isolated mathematical curiosity; it is an infinite-dimensional realization of Exact WKB analysis, Picard-Lefschetz resurgence, and the geometry of Spectral Networks [12, 8]. The complexified path integral over the boundary twistors exhibits localization phenomena that strictly mirror the topological mechanisms governing Seiberg-Witten (SW) theory and the analytic continuation of gauge theories [55]. To establish this physical correspondence, we analyze the three fundamental geometric mechanisms driving the twistor string quantization:
16.1 Topological Barriers and Infinite-Dimensional Picard-Lefschetz Theory
By analytically continuing the twistor path integral into the complexified phase space, the convergence of the quantum amplitude is governed by middle-dimensional integration cycles, or Lefschetz thimbles [4]. As formulated by Witten in the analytic continuation of Chern-Simons theory [55], the topology of these thimbles—and their intersection pairings—changes strictly when the system crosses a Stokes line, resulting in a discontinuous jump in the physical vacuum state. In our twistor string framework, non-perturbative Stokes phenomena are rigorously suppressed by the conformal boundary geometry. If a twistor pole $\xi_{1}$ attempts to cross the unit circle $|z|=1$, it inevitably collides with its conjugate mirror $\xi_{2}=1/\overline{\xi}_{1}$ directly on the physical integration contour. Exactly at the moment of crossing, these two singularities collide directly on the physical integration contour $|z|=1$. As demonstrated by the exact bulk-to-boundary reduction of the Liouville action in E, this collision creates a classic, non-integrable pinch singularity. The integration loop becomes trapped between the colliding singularities at $z_{0}$ and $1/\bar{z}_{0}$ and cannot be deformed to avoid them. This generates a severe pinch singularity, creating an infinite action barrier. In the language of Picard-Lefschetz theory, this impenetrable boundary dynamically locks the Twistor String path integral into a strictly conserved topological sector defined by the enclosed twistor pole number $n$, forbidding unphysical Stokes jumps.
16.2 Complex Langevin, Cauchy’s Shield, and Flat Valleys
The rigorous exploration of these complex integration cycles is executed via the holomorphic gradient drift of the Complex Langevin Equation (CLE). In standard semiclassical formulations, evaluating the path integral at a degenerate saddle point requires formally isolating a zero mode and calculating the fluctuation determinant of the Hessian. In the CLE framework, this static abstraction is bypassed entirely: the path integral expectation value is simply the long-time average $\int d\tau$ over a single, noisy stochastic trajectory $\mathcal{V}(\tau)$.
At the critical resonance energies $E=E_{n}$, the twistor effective action develops a non-compact flat valley. The single CLE trajectory does not rest at a classical minimum; driven by the stochastic noise, it constantly climbs the steep transverse walls of the complexified functional space and bounces back. Simultaneously, experiencing no deterministic restoring force along the zero mode, the trajectory drifts unimpeded along the flat valley. In this representation, the macroscopic pre-exponential factor (the transition residue) is simply proportional to the total Langevin time the trajectory spends accumulating windings, completely eliminating the need for formal Hessian factorization.
A fundamental mathematical question arises: how can this violently stochastic trajectory $\mathcal{V}(\tau)$ generate a pristine, exact Bohr-Sommerfeld quantization phase without acquiring infinite quantum corrections from the nonlinear effects amplifying the noise in transverse directions?
The resolution lies in the analytic continuation of the twistor fields. The CLE directly evolves the boundary data $\mathcal{V}_{bound}(\tau)=\mathcal{V}(z,\tau)\big|_{|z|=1}$. However, the twistors defining our symplectic form $\mathcal{O}$ are analytically continued from this boundary into the bulk. If we represent the twistor field as a meromorphic function, such as $\lambda(z)=\sum_{i}\frac{R_{i}}{z-z_{i}}$, its boundary values are:
$\lambda(\theta)=\sum_{i}\frac{R_{i}}{e^{i\theta}-z_{i}}$ (16.1)
The stochastic CLE evolution simply forces these boundary parameters to evolve with Langevin time $\tau$:
$\lambda(\theta,\tau)=\sum_{i}\frac{R_{i}(\tau)}{e^{i\theta}-z_{i}(\tau)}$ (16.2)
Upon analytic continuation back into the bulk, we obtain:
$\lambda(z,\tau)=\sum_{i}\frac{R_{i}(\tau)}{z-z_{i}(\tau)}$ (16.3)
The crucial observation is that the function remains strictly meromorphic at every instant of Langevin time. We obtain moving poles with evolving residues, but the fundamental meromorphic structure is absolutely preserved. Therefore, the infinite-dimensional stochastic evolution of the boundary fields rigorously and exactly reduces to the finite-dimensional, continuous $\tau$-dependence of the moduli $\{R_{i}(\tau),z_{i}(\tau)\}$.
The infinite-dimensional phase accumulation thereby rigidly reduces to a standard one-dimensional contour integral over the physical boundary:
$\Delta S=\oint_{|z|=1}\omega(z,\tau)\,dz$ (16.4)
Here, the absolute geometric power of Cauchy’s integral theorem manifests itself. At any instant $\tau$, this contour integral is completely determined by the sum of the residues at the enclosed twistor poles.
To evaluate the full physical amplitude, the path integral must be decomposed into a discrete sum over distinct topological sectors. These sectors are labeled by the winding number $k$ of the target-space minimal surface (or equivalently, the number of times the physical boundary parameter traverses the twistor poles). The full quantum amplitude is thus a superposition of path integrals, each mathematically explored by its own Complex Langevin evolution. Because the twistor effective action is strictly additive over these cycles, the total action for the $k$-th topological sector is precisely $S_{k}=S_{eff}^{(1)}+(k-1)\Delta S_{eff}$.
In the CLE formulation for the $k$-th sector, the stochastic noise is opposed by the deterministic drift and the transverse Hessian, which now scale linearly with the topological winding number: $\mathcal{H}_{k}\to k\nabla^{2}\Delta S_{eff}$. Consequently, the residual stochastic variance in this sector is inversely proportional to the Hessian, scaling exactly as $\langle\delta\mathcal{V}^{2}\rangle\propto 1/k$. If the CLE noise were to continuously jitter the moduli $\{R_{i},z_{i}\}$ across the different sectors, the topological action increment $\Delta S_{eff}$ would randomly vary from one winding to the next. This would destroy the perfect phase coherence required for the geometric series to diverge, preventing the formation of a resonance.
However, the macroscopic S-matrix pole is analytically generated exclusively by the asymptotic tail of this superposition, where $k\to\infty$. In this asymptotic limit, the transverse restoring force becomes strictly infinite, and the stochastic variance vanishes identically. The stochastic noise is completely crushed by the macroscopic weight of the multiple windings. The moduli $\{R_{i},z_{i}\}$ rigidly freeze at their exact classical saddle-point values for all large-$k$ sectors (validating the algebraic minimization of the single-pole radius $R$ in our spectral derivation). The topological phase is thus completely shielded from quantum contamination, ensuring that $\Delta S_{eff}$ remains strictly constant. The superposition over topological sectors safely diverges as a pristine geometric series $\sim(1-\exp\left(i\Delta S_{eff}\right))^{-1}$, analytically generating the exact S-matrix pole.
17 Conclusion: The Solution of Planar QCD

### 23 Mar 2026 v4

@@ 16.8 The saddle point equation
Crucially, the Fourier projection integral over the twist angle $\alpha$ inherently generates a kinematic factor of $1/k$:
$\int_{0}^{2\pi}d\alpha\,e^{ik\alpha\Phi_{\star}}=\frac{e^{2\pi ik\Phi_{\star}}-1}{ik\Phi_{\star}}\propto\frac{1}{k\Phi_{\star}}$ (16.35)
Combining this $1/k$ kinematic factor with the $k^{1/4}$ fluctuation weight, the terms in the macroscopic winding sum scale as $k^{1/4}/k=k^{-3/4}$. Summing this up to a macroscopic cutoff $k_{\text{max}}$ makes the divergence algebraically stronger than a logarithm, but preserves the fundamental geometric existence of the simple mass pole:
$A_{J}\propto\frac{1}{\Phi_{\star}(E,J)}\sum_{k=1}^{k_{\text{max}}}k^{-\frac{3}{4}}\propto\frac{(k_{\text{max}})^{\frac{1}{4}}}{\Phi_{\star}(E,J)}$ (16.36)
Strictly speaking, the integral remains finite at finite $k_{\text{max}}$ when $\Phi_{\star}(E,J)\to 0$. However, this happens in the infinitesimal vicinity, when $\Phi_{\star}(E,J)\sim 1/k_{\text{max}}$. Therefore, with proper renormalization of the string path integral, we can tend $k_{\text{max}}\to\infty$, after which we find the desired pole at $\Phi_{\star}(E,J)=0$ .
16.10 Comparison with Experimental Meson Spectrum
In this section, we compare the planar Regge trajectories derived from the twistor string effective action with the experimental meson mass spectrum from the Particle Data Group (PDG). The bound states are governed by the unified parametric equations derived in E. Parameterized by the spatial boundary phase $\beta\in[0,\pi/2)$ and the dimensionless renormalized boundary mass $x=m/\sqrt{\sigma}$, the energy $E=M$ and angular momentum $J$ of the states are (see Remark 16.2 for the definition of the renormalized boundary mass):
$\frac{M(\beta)}{\sqrt{\sigma}}=\mathcal{K}(\beta,x)\left(\beta+\frac{\sin 2\beta}{2}\right)+2x\cos\beta,$
$J(\beta)=\frac{\mathcal{K}(\beta,x)^{2}}{4}\left(\beta+\frac{\sin 2\beta}{2}\right)-\frac{1}{6\pi}(\tan\beta-\beta)-\frac{q}{2},$
@@ 16.10 Comparison with Experimental Meson Spectrum
$\mathcal{K}(\beta,x)=\frac{\sin\beta}{\cos^{2}\beta}\left(x+\sqrt{x^{2}-\frac{1}{3\pi}}\right)\equiv\frac{\sin\beta}{\cos^{2}\beta}\tilde{x}$. (16.37)
Here, the topological trace phase $q=0$ identifies the unshifted unnatural parity ($\pi,K,D$) trajectories, while q=1q=-1 shifts the natural parity ($\rho,K^{*},D^{*}$) trajectories by +1/2+1/2. To simplify the asymptotic expansions, we have introduced the dynamically dressed dimensionless mass parameter $\tilde{x}$, which enforces the chiral existence bound $x\geq 1/\sqrt{3\pi}$.
Remark 16.2.
The Renormalized Boundary Mass. The chiral existence bound $x\geq 1/\sqrt{3\pi}$ forbids the limit of a massless string boundary. The geometric drag generates a minimum inertial mass to balance the anomaly at the boundaries. The parameter $m$ appearing in the string effective action is not the bare QCD quark mass, nor is it the standard phenomenological constituent mass. It is a renormalized boundary mass. It represents the bare quark mass additively renormalized by the macroscopic perimeter term of the worldsheet fermion determinant, inducing the Planar QCD in our theory.
Remark 16.3.
Choosing the lightest particles (ground state). In general, there are several resonances in the PDG tables with the same quantum numbers $I^{G}(J^{C})$ in each meson sector (given strangeness $S$ and charm $C$). We always selected the lowest mass particle with a constant set of quantum numbers at given variable $J$ to place on the same Regge trajectory $J=\alpha(M^{2})$. Presumably, higher resonances in PDG tables can also be explained by twistor monodromies, but with more than one branching point. We leave this problem for a future extended investigation.
Universal String Tension and the Global Fit. Experimentally, the asymptotic slopes of the meson trajectories differ by less than $10\%$. Within the planar limit ($N_{c}\to\infty$) of Geometric QCD, there exists only one universal gluon vacuum and a single, universal confining string tension $\sigma$. At leading order in $1/N_{c}$, there is no mechanism to produce different macroscopic slopes for different flavors. The observed $10\%$ splitting is an $\mathcal{O}(1/N_{c}^{2})$ non-planar effect driven by finite resonance widths, sea-quark loop insertions, and string-string interactions. This falls within the expected theoretical accuracy of the planar approximation.
Consequently, we perform a global joint optimization across the thirteen meson families ($\pi,\rho,K,K^{*},D,D^{*},D_{s},D_{s}^{*},B,B^{*},B_{s},B_{s}^{*},B_{c}$). Instead of an arbitrary least-squares fit on masses, we minimize the fundamental action monodromy $\sum(\Delta S)^{2}$ across all thirty-six states. This global 5D optimization extracts exactly five fundamental parameters: the universal planar string tension $\sqrt{\sigma}$, and the four constituent boundary masses of the quarks ($m_{u},m_{s},m_{c},m_{b}$). The asymmetric heavy-light and heavy-heavy trajectories mathematically utilize the arithmetic mean of their respective constituent boundary masses.
The global fit yields a universal string tension $\sqrt{\sigma}\approx 417$ MeV and constituent boundary masses $m_{u}\approx 136$ MeV, $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV. The fact that the $D_{s}$, $B_{s}$, and $B_{c}$ trajectories are completely locked by the relation $\bar{m}_{q_{1}q_{2}}=(m_{q_{1}}+m_{q_{2}})/2$ without introducing any new free parameters proves that the boundary masses physically correspond to additive constituent quarks in the string endpoints. The ability of a single geometric tension to seamlessly govern the spectrum from the $140$ MeV pion all the way up to the $6.27$ GeV bottom resonances validates the absolute geometric universality of the twistor string.
The resulting non-linear trajectories are presented against the PDG data in Figure 15.
Figure 15: The geometric Regge trajectories plotted against the PDG meson spectrum. A single universal planar string tension $\sqrt{\sigma}\approx 417$ MeV anchors all 13 families. The $\pi$ and $\rho$ trajectories generate an $m_{u}\approx 136$ MeV renormalized boundary mass. The heavier trajectories accommodate the strange, charm, and bottom quarks via the arithmetic means of their dynamically extracted constituent boundary masses: $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV. The topological spin shift q=1q=-1 elevates the natural parity branches. Experimental deviations in local slopes ($\sim 10\%$) are consistent with expected $\mathcal{O}(1/N_{c}^{2})$ non-planar corrections.
The Ultra-Relativistic Asymptote (Large $J$). To analytically evaluate the high-energy limit, we expand the parametric equations as the boundaries approach the speed of light ($\beta\to\pi/2$). Defining the small parameter $\epsilon=\pi/2-\beta\to 0$, the trigonometric functions expand as $\cos\beta\approx\epsilon$ and $\sin\beta\approx 1-\epsilon^{2}/2$. The geometric equations expand to leading order as:
$E(\epsilon)$ $\approx\frac{\pi\tilde{x}\sqrt{\sigma}}{2\epsilon^{2}}\implies\frac{1}{\epsilon}\approx\sqrt{\frac{2E}{\pi\tilde{x}\sqrt{\sigma}}},$
$J(\epsilon)$ $\approx\frac{\pi\tilde{x}^{2}}{8\epsilon^{4}}-\frac{\pi\tilde{x}^{2}}{24\epsilon^{2}}-\frac{1}{\epsilon}\left(\frac{\tilde{x}^{2}}{6}+\frac{1}{6\pi}\right)+\frac{1}{12}-\frac{q}{2}.$ (16.38)
Because the squared energy expands as $\frac{E^{2}}{2\pi\sigma}\approx\frac{\pi\tilde{x}^{2}}{8\epsilon^{4}}-\frac{\pi\tilde{x}^{2}}{24\epsilon^{2}}+\frac{x\tilde{x}-\tilde{x}^{2}/3}{\epsilon}$, substituting the inverted $1/\epsilon$ relation into the angular momentum equation cancels the divergent extensive terms, revealing the high-energy Regge trajectory:
$J(E)\approx\frac{E^{2}}{2\pi\sigma}-\left(x\tilde{x}-\frac{\tilde{x}^{2}}{6}+\frac{1}{6\pi}\right)\sqrt{\frac{2E}{\pi\tilde{x}\sqrt{\sigma}}}+\frac{1}{12}-\frac{q}{2}.$ (16.39)
This expansion modifies the linear Nambu-Goto expectation. The constant conformal intercept stabilizes to +1/12+1/12, recovering the Lüscher zero-point energy of the open bosonic string in $D=4$. However, the Liouville anomaly mandates a finite effective boundary mass $\tilde{x}$, which acts as a geometric drag, generating a negative $\mathcal{O}(\sqrt{E})$ curvature. At large $J$, relativistic effects act on the massive endpoints, and this geometric inertial lag curves the trajectories downward. This departs from the straight lines of Nambu-Goto models and reproduces the slope-softening observed in physical high-spin resonances.
The Non-Relativistic Hook (Small $J$). While the high-energy limit is dominated by relativistic drag, the finite renormalized boundary mass $m=x\sqrt{\sigma}$ generates a geometric non-linearity confined to the low-$J$ region. Near $J\to 0$, the string is short, the endpoints move non-relativistically, and the boundary angle shrinks ($\beta\to 0$). Taylor-expanding the parametric equations around $\beta=0$ yields:
$E(\beta)$ $\approx 2m+\beta^{2}\sqrt{\sigma}\left(2\tilde{x}-x\right),$
$J(\beta)$ $\approx\beta^{3}\left(\frac{\tilde{x}^{2}}{2}-\frac{1}{18\pi}\right)-\frac{q}{2}.$ (16.40)
Because the kinetic energy scales quadratically ($\Delta E\propto\beta^{2}$) while the angular momentum scales cubically ($J\propto\beta^{3}$), isolating $\beta$ from the energy equation and substituting it into $J$ produces the non-relativistic trajectory:
$J(E)\approx\left(\frac{\tilde{x}^{2}}{2}-\frac{1}{18\pi}\right)\left(\frac{E-2m}{\sqrt{\sigma}(2\tilde{x}-x)}\right)^{3/2}-\frac{q}{2}.$ (16.41)
Due to the fractional $3/2$ power, the macroscopic slope of the Regge trajectory at the threshold evaluates to $dJ/dE\propto\sqrt{E-2m}\to 0$. This demonstrates that the trajectories form a geometric hook, entering the physical parameter space with a flat derivative before accelerating into the linear regime. This non-relativistic hook is beautifully realized in the heavy charmed ($D,D^{*}$) and bottom ($B,B_{s},B_{c}$) meson trajectories, which sit deep in the non-relativistic regime for low spins before curving upward towards the universal linear asymptote.
Phenomenological Synthesis. The global geometric optimization of the dimensionless action monodromy target ($\sum(\Delta S)^{2}=0$) fits the experimental PDG masses and spins across all thirteen meson families. By evaluating the parametric equations of the twistor string rather than linear approximations, several phenomenological features of QCD are recovered.
First, the trajectories demonstrate the dynamical generation of a renormalized boundary mass gap. Unlike classical Nambu-Goto strings, which yield linear trajectories passing through the origin, the integration of the Liouville anomaly bounds the dimensionless mass to $x\geq 1/\sqrt{3\pi}$, forbidding the massless string limit. The theoretical trajectories terminate at finite rest masses. The geometric pressure of the anomaly forces the string endpoints to generate a renormalized boundary mass of $m_{u}\approx 136$ MeV for the light sector. For the heavier sectors, the 5D optimization dynamically extracts the constituent boundary masses $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV. The asymmetric heavy trajectories mathematically utilize the arithmetic mean of these respective scales, naturally isolating the physical flavor-symmetry breaking scales across the entire Standard Model meson spectrum.
Second, the topological intercept distinguishes the vector from the pseudo-scalar mesons. Setting the topological vector index q=1q=-1 for the for the vector families ($\rho,K^{*},D^{*},D_{s}^{*},B^{*},B_{s}^{*}$) shifts their respective trajectories upward without arbitrary additive constants. The zero-point rotational quantum phase converts the generated constituent rest mass into the necessary kinetic energy, predicting the physical $J=1$ vector ground states.
Third, the high-energy asymptotic regime reveals the effect of the conformal anomaly. The boundary anomaly term $\frac{1}{6\pi}(\tan\beta-\beta)$ behaves as a geometric drag. As the massive string endpoints approach the speed of light ($\beta\to\pi/2$), this inertial lag curves the trajectories downward, reproducing the slope-softening observed in physical resonances ($M^{2}>4\text{ GeV}^{2}$), departing from the straight lines of Nambu-Goto models.
Finally, the systematic variations of the experimental data points around the theoretical curves demonstrate the model’s assumption of a universal underlying string tension $\sigma$. In physical QCD, explicit chiral symmetry breaking and hyperfine spin interactions cause the effective Regge slopes to vary slightly between different families. By enforcing a universal vacuum tension, the theoretical target $\sum(\Delta S)^{2}=0$ balances this physical variance. The resulting curves represent a universal tension of the strong vacuum that connects the distinct features of the hadronic spectrum.
Numerical Mass Predictions and Fit Accuracy.
State $J$ $M_{\rm th}$ (GeV) $M_{\rm exp}$ (GeV) State $J$ $M_{\rm th}$ (GeV) $M_{\rm exp}$ (GeV)
Light Pseudoscalars ($\pi$, $q=0$) Light Vectors ($\rho$, q=1q=-1)
$\pi$ 0 0.271 0.140 $\rho$ 1 0.878 0.775
$b_{1}$ 1 1.168 1.229 $a_{2}$ 2 1.394 1.318
$\pi_{2}$ 2 1.586 1.671 $\rho_{3}$ 3 1.756 1.689
$a_{4}$ 4 2.052 1.967
$\rho_{5}$ 5 2.308 2.330
Strange Pseudoscalars ($K$, $q=0$) Strange Vectors ($K^{*}$, q=1q=-1)
$K$ 0 0.355 0.494 $K^{*}$ 1 0.928 0.892
$K_{1}$ 1 1.212 1.272 $K^{*}_{2}$ 2 1.435 1.430
$K_{2}$ 2 1.624 1.580 $K^{*}_{3}$ 3 1.792 1.780
$K^{*}_{4}$ 4 2.085 2.045
$K^{*}_{5}$ 5 2.340 2.380
Charm Pseudoscalars ($D$, $q=0$) Charm Vectors ($D^{*}$, q=1q=-1)
$D$ 0 1.739 1.865 $D^{*}$ 1 2.118 2.008
$D_{1}$ 1 2.331 2.422 $D^{*}_{2}$ 2 2.506 2.463
$D_{2}$ 2 2.660 2.747 $D^{*}_{3}$ 3 2.799 2.763
Charm-Strange ($D_{s}$, $q=0$) Charm-Strange Vectors ($D_{s}^{*}$, q=1q=-1)
$D_{s}$ 0 1.823 1.968 $D_{s}^{*}$ 1 2.196 2.112
$D_{s1}$ 1 2.407 2.460 $D_{s2}^{*}$ 2 2.580 2.569
$D_{s3}^{*}$ 3 2.869 2.860
Bottom Pseudoscalars ($B$, $q=0$) Bottom Vectors ($B^{*}$, q=1q=-1)
$B$ 0 5.207 5.279 $B^{*}$ 1 5.475 5.325
$B_{1}$ 1 5.630 5.726 $B_{2}^{*}$ 2 5.760 5.739
Bottom-Strange ($B_{s}$, $q=0$) Bottom-Strange Vectors ($B_{s}^{*}$, q=1q=-1)
$B_{s}$ 0 5.291 5.367 $B_{s}^{*}$ 1 5.557 5.415
$B_{s1}$ 1 5.711 5.829 $B_{s2}^{*}$ 2 5.840 5.840
Charm-Bottom ($B_{c}$, $q=0$)
$B_{c}$ 0 6.675 6.275
Table 1: The theoretical twistor string masses versus the experimental PDG meson masses. All theoretical values are generated from a single global $\sum(\Delta S)^{2}=0$ optimization over 36 states spanning 13 meson families, extracting exactly 5 fundamental physical parameters: a universal planar tension $\sqrt{\sigma}\approx 417$ MeV, and four dynamically generated constituent boundary masses $m_{u}\approx 136$ MeV, $m_{s}\approx 219$ MeV, $m_{c}\approx 1.60$ GeV, and $m_{b}\approx 5.07$ GeV. The asymmetric heavy trajectories use the arithmetic mean of the respective constituent boundary masses. Therefore, the $D_{s},B_{s},B_{c}$ families are parameter-free geometric predictions. The pion falls below the string’s mass threshold ($2m_{u}\approx 271$ MeV) due to explicit chiral symmetry breaking.
To quantify the accuracy of the planar geometric fit, rather than relying on artificial least-squares statistics, we evaluate three physical root-mean-square (RMS) metrics across the experimental spectrum. The theoretical mass predictions for all thirty-six on-shell states are presented in Table 1.
The most fundamental geometric metric for the fit is the residual of the action monodromy phase $\Delta\Phi=\Delta S/a$. In the Bohr-Sommerfeld quantization condition, sequential discrete states are separated by a phase gap of exactly $\Delta S_{\rm gap}/a=2\pi$. Evaluating the Root-Mean-Square (RMS) of the action residual across all 36 states yields ${\rm RMS}(\Delta\Phi/2\pi)\approx 0.36$. This demonstrates that the geometric phase error is merely $\sim 1/3$ of a single quantum level, ensuring that the physical states remain safely isolated within the $n=0$ topological winding sector without overlapping into radial excitations.
Furthermore, because the non-linear Regge trajectories are defined geometrically as $J=\alpha(M^{2})$, a natural physical metric for their accuracy is the absolute spin deviation. Evaluating the theoretical angular momentum at the exact experimental masses yields an RMS deviation of $\Delta J_{\rm RMS}\equiv\sqrt{\langle(J_{\rm exp}-\alpha(M_{\rm exp}^{2}))^{2}\rangle}\approx 0.18$. Since physical Regge states are separated by exactly $\Delta J=1$, this confirms the twistor geometry correctly threads the mass spectrum to within less than one-fifth of an orbital quantum step.
Finally, we evaluate the relative mass deviations. Excluding the $140$ MeV pion (which acts as an explicit pseudo-Goldstone boson and sits naturally below the geometric chiral threshold $2m_{u}\approx 271$ MeV), the RMS of the relative mass error $(M_{\rm exp}/M_{\rm theor}-1)$ across the remaining 35 massive states is $\approx 8.1\%$. This firmly aligns with the expected $\mathcal{O}(1/N_{c}^{2}\approx 10\%)$ theoretical variance of planar dynamics at $N_{c}=3$, verifying that the universal planar string tension accurately anchors the disparate bottom, charm, strange, and light flavors.
17 Conclusion: A Solution of Planar QCD